# Optimal market-neutral currency trading on the cryptocurrency platform

### v2

@@ 5.1 Experimental Setup
We trade a bucket of fiat currency pairs referenced to ETH for a number of reasons: ① Fiat currency pairs (equivalently spreads) referenced to ETH are cointegrated and stationary; ② Cryptocurrency markets have substantial liquidity and depth; ③ The volatility in the cryptocurrency market overwhelms the volatility inherent volatility in the fiat, which gives us opportunities to arbitrage. We adopt the trading bucket in the ETH pairs on the Kraken [^1] https://www.kraken.com/ exchange, where the assets are freely accessible to all participants and yet traded independently. We select the cryptocurrency exchange Kraken because it is one of the top exchanges providing most fiat options for platform users without geographical restrictions. Following this practice, we are adopting mainstream fiat currencies with considerable market activity toward cryptocurrency and have been traded for a substantial period of time.
The datasets we use are publicly available from Kraken OHLCVT dataset [^2] The dataset is downloadable at: https://support.kraken.com/hc/en-us/articles/360047124832-Downloadable-historical-OHLCVT-Open-High-Low-Close-Volume-Trades-data, which is the historical trading data from Kraken exchange for Open price, Highest price, Lowest price, Close price, trading Volume and number of Trades at different sampling time intervals. The sampling time intervals adopted in this experiment are 1 min, 5 min and 60 min.
@@ 5.1.1 Accounting for transaction fees
Transaction cost in quantitative trading is always a non-negligible factor to be considered. One of the incentives to introduce optimal trading is the benefit of reducing transaction costs. Buy-and-hold strategy receives the least impact from transaction costs.
Kraken exchange sets a volume-based fee scheme. The more trading volume, the lower the transaction fee can be. In practice, quantitative trading normally involves large trading volumes. We adopt a flat 0.1% transaction fee for long and short orders.
@@ 5.2 Data Source Screening
As mentioned previously (c.f. Section 3.1), a preliminary requirement of the proposed OTT is that fiats quoted for ETH are correlated and cointegrated. If not, it indicates that trading on the bucket of Fiat/ETH pair is hardly following a mean-reversion process. We use the SPOT ETH price quoted in different USD, CAD, GBP, and EUR from 01-Jan-2018 to 01-Jan-2020 for screening and formation (Algorithm 1 line 1). Our dataset is sampled at 1min, 5min and 60min trading intervals. We want to test the stationarity of the pairs (spreads) not only for the complete formation period, but also for shorter cycles corresponding to the time frames of our arbitrage. A low correlation or cointegration result implies bad applicability of the proposed OTT. Therefore we examine the data on a rolling 1-week basis to test for correlation and stationarity via cointegration.
Table 2: Correlation and Cointegration Among Currency Pairs
@@
  • This table provides average rolling correlation calculation and cointegration test results for each currency pair between 01-Jan-2018 and 01-Jan-2020.
The correlation and cointegration test results are presented in Table 2. USD:CAD under 1-week stands for Pearson’s correlation for USD:CAD between 01-Jan-2018 to 01-Jan-2020 cut into 105 weeks. The correlation we present is the simple weekly average for 105 weeks. We expect our trading target to be stationary. Therefore we also expect the spreads we trade to be cointegrated during the period we hold the positions. Similar to Pearson’s correlation test, the data are also cut into weekly slots for the Engle-Granger test. The passing percentages are presented in the table. We can see that the data generally have good correlations and cointegration relationships, so we can confidently conduct the proposed arbitrage.
@@ 5.3 Tuning Trading Parameters
Once the stationarity of the spreads is determined we compute a number of parameters including thresholds, which determine the boundaries to open and close position from the formation data (Algorithm 1 line 1).
@@
The two parameters that determine the closing and opening boundaries are Open_Threshold and Close_Threshold multiplied by the standard deviation (c.f. Table 1). Given that the assets are naturally different, and we are betting on the volatility of spreads, therefore we devise parameters based on the standard deviation. We must decide carefully on what values will be used for the trading signals. In our research, we conduct grid-search for the thresholds that are likely to generate high profitability. Figure 5 illustrates the grid search results for 5min sample interval formation dataset. It is a heatmap for the indicators of profitability (calculation method in Section 5.3.1). A darker shade means the better profit of the Open_Threshold/Close_Threshold combination. We can see that in our test phase, for 5min sampled dataset, opening at 9 standard deviations and closing at 7 standard deviations generates the best profitability. The grid search is extended to the trading intervals of 5min and 60min as well. Table 3 records the most lucrative combination for adopted trading intervals in this research.
Table 3: The most profitable trading threshold
@@
  • The thresholds are calculated from retrospective grid searches from 05-Jan-2020 and 01-Oct-2022 for the most lucrative combination of Open_Threshold and Close_Threshold.
Another tunable parameter is the risk-preference parameter LAMBDA ($\lambda$), which is used to adjust the punishment factor for investment risk in Equation (2). We can tweak $\lambda$ to cater for different risk preferences. A high $\lambda$ represents a strong punishing factor for portfolio volatility, leading to a more conservative investment decision. By default, $\lambda$ is set to 1. But we will also examine the profitability of arbitrage under a riskier $\lambda$=0.5 and a more conservative $\lambda$=2.
@@ 5.3.1 Calculating P&L in base currency
Since we are trading with multiple fiat currencies, we need a unified method to measure profitability (or loss) — so-called P&L. The way we do this is to choose a base currency among participating assets — USD in our case. We obtain the foreign currency exchange rate for USD-CAD, USD-EUR, and USD-GBP from the beginning of the arbitrage testing period to the end (Table 4).
Table 4: Foreign Currency Exchange Rate for USD
Full-Cycle Market 15-Jan-2021 01-Oct-2022
USD : USD 1.000000 1.000000
USD : CAD 1.266453 1.383270
USD : GBP 0.758367 0.896861
USD : EUR 0.888224 1.019910
Bull Market 01-Jan-2021 01-Jan-2022
USD : USD 1.000000 1.000000
USD : CAD 1.273005 1.263930
USD : GBP 0.731368 0.739196
USD : EUR 0.821600 0.879569
Bear Market 01-Jan-2022 01-Jan-2023
USD : USD 1.000000 1.000000
USD : CAD 1.263930 1.354530
USD : GBP 0.739196 0.826327
USD : EUR 0.879569 0.931797
Notice that our testing dataset does not overlap with the formation period (12-Mar-2016 — 31-Dec-2021). Hence, all our arbitrage tests are on out-of-sample (never seen before) datasets.
We start with 10000 USD, CAD, GBP and EUR, each. It is equivalent to holding 42340.72 USD on 15-Jan-2021. If all the assets we hold remain the same, it will be converted as equivalent to holding 38184.03 USD on 01-Oct-2022. The assets we kept are the same, but we calculated the result as a 9.82% loss. However, it does not mean that we will naturally lose 9.82% due to no action. Since our technique tracks forex volatility as part of the trading signal, a bullish USD will navigate the arbitrage to long ETH with USD for other currencies. So we can view our profitability as a compounded effect of cryptocurrency arbitrage and forex volatility arbitrage. Another example with portfolio changes is presented in Table 5. The starting portfolio is converted to USD as stated above. The end date portfolio is converted based on end date FX. Hence we can have a comparable profitability of 42340.72 USD to 40564.62 USD as -4.19%.
@@ 5.4.1 Full-Cycle market
We use the method in Section 5.3.1 to calculate the profitability in USD, presented in Table 6. The numbers in the table for various currencies are the profit (or loss) at the end of the trading period, starting with 10000 amount each. All currencies are converted to USD based on the exchange rate at the start of the trading period of 15-Jan-2021. We convert the currencies back to USD for profitability calculation on 01-Oct-2022. During this period, cryptocurrency experienced a bull market and a bear market, allowing us to test our technique through peak and trough. The results of our technique are encouraging. We achieved an annualized return of 14.87% in USD for 1min sampled dataset ($\lambda$=1). For the 5min and 60min sampled datasets, the annualized returns were 15.49% and 5.00%, which are greatly exceeding the Distance Method (DM) in all experimented cases.
Figure 6: The profitability of Full-Cycle Market between 15-Jan-2021 to 01-Oct-2022
We also conduct experiments on the conditions where investors have higher or less tolerance for portfolio volatility. We can observe that there is a clear relationship: the lower the $\lambda$, the higher the profitability. This relationship meets our expectation that the investors receive a risk premium in exchange for a more volatile portfolio.
Table 6: Annualized profitability results for $\lambda$ in full-cycle market (Adjusted)
Interval BH DM OTT($\lambda$=0.5) OTT($\lambda$=1) OTT($\lambda$=2)
1 min (TC 0.1%) 13.14% 5.33% 35.54% 14.87% 7.44%
1 min (TC 0%) 13.14% 0.73% 79.98% 40.03% 19.97%
5 min (TC 0.1%) 13.14% 0.60% 37.74% 15.49% 7.74%
5 min (TC 0%) 13.14% 4.08% 90.03% 44.98% 22.49%
60 min (TC 0.1%) 13.14% 3.93% 9.93% 5.00% 2.51%
60 min (TC 0%) 13.14% 5.54% 37.43% 18.96% 9.49%
  • 1
    The arbitrage period is 15-Jan-2021 to 01-Oct-2022. Participating assets are USD, CAD, GBP, and EUR. The absolute return is stated in percentages.
  • 2
    Abbreviations. BH=”Buy-Hold”, DM=”Distance Method”, TC=”Transaction Cost”, OTT=”Optimal Trading Technique”
  • 3
    Open_Threshold and Close_Threshold are tuned best on the most lucrative combination (Table 3)
Another comparison is the inclusion of transaction costs. We can see that even a 0.1% transaction cost significantly impacts our returns. On the surface, transaction cost charges a flat percentage of trading capital and influences the optimization process’s expectation. Our technique trades on a frequent basis; hence we can see the profitability is almost halved in three examined $\lambda$ values after including transaction costs. The Distance Method, on the other hand, does not have a significant impact due to transaction cost, because it trades less frequently (Table 10).
@@ 5.4.2 Bull market
Figure 7: The profitability of Bull Market between 01-Jan-2021 to 01-Jan-2022
Table 7 presents the trading results for 2021. A significant growth trend is evident from the Buy-Hold strategy, which yielded a return of 412.09%. Such a circumstance is rare and occurs only in exceptional cases. Compared to full-cycle market trading, both the Distance Method (DM) and our Optimal Trading Technique (OTT)yielded more profit in the bull market. However, it was still lower than the simple Buy-Hold strategy. For a standard scenario of $\lambda=1$ and $tc=0.1\%$ over a five-minute period, the Optimal Trading Technique achieved an annual return of 36.75% (compared to 19.89% for the Distance Method under the same conditions). Single-sided markets are easier to estimate, and the expected_return and expected_risk are therefore closer to reality. The profitability trend is consistent with that of the full-cycle market. Arbitrages that do not entail transaction fees generate better returns than those with transaction fees. A lower value of Lambda ($\lambda$) implies a higher acceptance risk, resulting in a better overall arbitrage return.
Table 7: Annualized profitability results for $\lambda$ tuning in bull market
Interval BH DM OTT($\lambda$=0.5) OTT($\lambda$=1) OTT($\lambda$=2)
@@
60 min (TC 0%) 412.90% 7.46% 45.25% 22.75% 11.37%
  • 1
    The arbitrage period is 01-Jan-2021 to 01-Jan-2022. Participating assets are USD, CAD, GBP, and EUR. The annualized return is stated in percentages.
  • 2
    Abbreviations. BH=”Buy-Hold”, DM=”Distance Method”, TC=”Transaction Cost”, OTT=”Optimal Trading Technique”
  • 3
    Open_Threshold and Close_Threshold are tuned best on the most lucrative combination (Table 3)
@@ 5.4.3 Bear market
We also have an interest in examining how well our technique performs in a bear market. The idea of pair trading is to pursue market neutrality, and our optimization process calculates weights to avoid including ETH in our portfolio. In theory, the process should perform as well as a bull market. Therefore, we experiment with the technique under a bear market year 2022 from 01-Jan-2022 to 31-Dec-2022. During this period, the price of an ETH dropped -61.77%. We examine the profitability with data for this price dip in Table 8.
Figure 8: The profitability of Bear Market between 01-Jan-2022 to 01-Jan-2023
During the bear market in 2022, ETH witnessed a significant drop in all major currencies. The traditional Buy–Hold strategy and most non-market-neutral techniques would experience a loss from the bear market. By not ETH at all, both Distance Method and our Optimal Trading Technique (OTT)achieved positive returns in all experimented cases. We can see from Table 8 that the patterns we found in a full-cycle market also apply here.
Table 8: Annualized profitability results for $\lambda$ tuning in bear market
Interval BH DM OTT($\lambda$=0.5) OTT($\lambda$=1) OTT($\lambda$=2)
@@
60 min (TC 0%) -61.77% 5.53% 33.08% 13.93% 8.52%
  • 1
    The arbitrage period is 01-Jan-2022 to 01-Jan-2023. Participating assets are USD, CAD, GBP, and EUR. The annualized return is stated in percentages.
  • 2
    Abbreviations. BH=”Buy-Hold”, DM=”Distance Method”, TC=”Transaction Cost”, OTT=”Optimal Trading Technique”
  • 3
    Open_Threshold and Close_Threshold are tuned best on the most lucrative combination (Table 3)

### v3

@@ 5.1 Experimental Setup
An overarching view of our method is to pair trade fiat currencies with a cryptocurrency as an intermediate. We adopt the trading bucket from Kraken [^2] https://www.kraken.com/ exchange, where assets trading against multiple fiat currencies are freely accessible to individual researchers. We select the cryptocurrency exchange Kraken because it is one of the top exchanges providing most fiat options for platform users without geographical restrictions.
We trade a bucket of fiat currency pairs referenced to ETH for several reasons: ① Cryptocurrency markets have substantial liquidity and depth; ② The volatility in the cryptocurrency market overwhelms the inherent volatility in fiat currencies, providing opportunities for arbitrage. ③ For fiat currency pairs (equivalently spreads) referenced to cryptocurrencies, ETH has the strongest cointegrated and stationary relationship comparing to BTC (c.f. Table 3, LABEL:tab:btc_corr);
Currencies with the most consistent trading volumes available on the Kraken exchange are USD, CAD, GBP, and EUR. The period from 15-Jan-2022 to 01-Oct-2022 represents a perfect full-cycle market where ETH prices against USD climb roughly quadruple in the first half and fall back to the beginning price in the second half. For more verifications, we also examined a whole-year bull market (01-Jan-2021 to 01-Jan-2022) and a whole-year bear market (01-Jan-2022 to 01-Jan-2023).
The datasets we use are publicly available from the Kraken OHLCVT dataset [^3] The dataset is downloadable at: https://support.kraken.com/hc/en-us/articles/360047124832-Downloadable-historical-OHLCVT-Open-High-Low-Close-Volume-Trades-data, which includes historical trading data from Kraken exchange for Open price, Highest price, Lowest price, Close price, trading Volume, and number of Trades at different sampling time intervals.
@@ 5.1.1 Accounting for transaction fees
Transaction costs in quantitative trading are always a non-negligible factor. One of the incentives to introduce optimal trading is the benefit of reducing transaction costs. The buy-and-hold strategy is least affected by transaction costs.
Kraken Exchange uses a volume-based fee scheme, where higher trading volumes result in lower transaction fees. In practice, quantitative trading typically involves large trading volumes. For our analysis, we adopt a flat 0.1% transaction fee for both long and short orders.
@@ 5.2 Data Source Screening
As mentioned previously (c.f. Section 3.1), a preliminary requirement of the proposed OTT is that fiat currencies quoted for ETH are correlated and cointegrated. If not, it indicates that trading on the bucket of Fiat/ETH pairs is unlikely to follow a mean-reversion process. We use the SPOT ETH price quoted in USD, CAD, GBP, and EUR from 01-Jan-2018 to 01-Jan-2020 for screening and formation (Algorithm 1 line 1). We test the stationarity of the pairs (spreads) not only for the complete formation period but also for shorter cycles corresponding to the time frames of our arbitrage. A low correlation or cointegration result implies poor applicability of the proposed OTT. Therefore, we examine the data on a rolling 1-week basis to test for correlation and stationarity via cointegration.
Table 2: ETH Correlation and Cointegration among Currency Pairs
@@
  • This table provides average rolling correlation calculations and cointegration test results for each currency pair between 01-Jan-2018 and 01-Jan-2020.
Table 3: ETH Correlation and Cointegration among Currency Pairs
Interval 1min Interval 5min Interval 60min
Currencies Corr Coint Corr Coint Corr Coint
USD:CAD 0.978 (77.8%) 0.998 (100.0%) 0.965 (89.9%)
USD:GBP 0.916 (66.7%) 0.974 (94.1%) 0.908 (72.5%)
USD:EUR 0.997 (77.8%) 0.914 (72.5%) 0.994 (69.6%)
CAD:GBP 0.917 (85.2%) 0.991 (72.5%) 0.903 (81.2%)
CAD:EUR 0.978 (88.9%) 0.980 (94.1%) 0.969 (98.6%)
GBP:EUR 0.917 (88.9%) 0.921 (98.0%) 0.913 (91.3%)
  • This table provides average rolling correlation calculations and cointegration test results for each currency pair between 01-Jan-2018 and 01-Jan-2020.
The correlation and cointegration test results are presented in Table 3. USD:CAD under 1-week represents Pearson’s correlation for USD:CAD between 01-Jan-2018 and 01-Jan-2020, divided into 105 weeks. The correlation presented is the simple weekly average for these 105 weeks. We expect our trading targets to be stationary, so we also expect the spreads we trade to be cointegrated during the period we hold the positions. Similar to the Pearson’s correlation test, the data are also divided into weekly slots for the Engle-Granger test. The passing percentages are presented in the table. We can see that the data generally have good correlations and cointegration relationships, so we can confidently conduct the proposed arbitrage.
@@ 5.3 Tuning Trading Parameters
Once the stationarity of the spreads is examined, we compute several parameters, including thresholds, which determine the boundaries to open and close positions based on the formation data (Algorithm 1 line 1).
@@
The two parameters that determine the closing and opening boundaries are open_threshold and close_threshold (c.f. Table 1). In our research, we conduct a grid search for the thresholds that are likely to generate high profitability. Figure 5 illustrates the grid search results for the 5min sample interval formation dataset. It is a heatmap for the indicators of profitability (calculation method in Section 5.3.1). A darker shade means better profit for the open_threshold/close_threshold combination. For the 5min sampled dataset, opening at 9 standard deviations and closing at 7 standard deviations generates the best profitability. The grid search is extended to the trading intervals of 1min and 60min as well. Table 4 records the most lucrative combination for the trading intervals adopted in this research.
Table 4: The most profitable trading thresholds
@@
  • The thresholds are calculated from retrospective grid searches from 05-Jan-2020 to 01-Oct-2022 for the most lucrative combination of open_threshold and close_threshold.
Another tunable parameter is the risk-preference parameter $\lambda$, which is used to adjust the penalty factor for investment risk in Equation (2). We can tweak $\lambda$ to cater to different risk preferences. A high $\lambda$ represents a strong penalty factor for portfolio volatility, leading to a more conservative investment decision. By default, $\lambda$ is set to 1, but we will also examine the profitability of arbitrage under a riskier $\lambda=0.5$ and a more conservative $\lambda=2$.
@@ 5.3.1 Calculating P&L in Base Currency
Since we are trading with multiple fiat currencies, we need a unified method to measure profitability (or loss)—so-called P&L. We choose a base currency among the participating assets, which is USD in our case. We obtain the foreign currency exchange rates for USD-CAD, USD-EUR, and USD-GBP from the beginning of the arbitrage testing period to the end. Note that our testing dataset does not overlap with the formation period (12-Mar-2016 to 31-Dec-2021). Hence, all our arbitrage tests are on out-of-sample (never seen before) datasets.
We start with 10,000 USD, CAD, GBP, and EUR each, equivalent to holding 42,340.72 USD on 15-Jan-2021. If all the assets we hold remain the same, they will be equivalent to holding 38,184.03 USD on 01-Oct-2022. Although the assets are unchanged, this calculation shows a 9.82% loss due to currency exchange rate changes. However, this does not imply a natural 9.82% loss without action. Since our technique tracks forex volatility as part of the trading signal, a bullish USD will navigate the arbitrage to long ETH with USD for other currencies. Thus, our profitability is a compounded effect of cryptocurrency arbitrage and forex volatility arbitrage. Another example with portfolio changes is presented in Table 5. The starting portfolio is converted to USD as stated above. The end-date portfolio is converted based on the end-date foreign exchange rate. Hence, we can compare profitability from 42,340.72 USD to 40,564.62 USD, resulting in a -4.19% change.
@@ 5.4.1 Full-Cycle Market
We use the method in Section 5.3.1 to calculate the profitability in USD, presented in Table 6. The numbers in the table for various currencies represent the profit (or loss) at the end of the trading period, starting with 10,000 units of each currency. All currencies are converted to USD based on the exchange rate at the start of the trading period on 15-Jan-2021. We convert the currencies back to USD for profitability calculation on 01-Oct-2022. During this period, cryptocurrency experienced both a bull run and a bear run, allowing us to test our technique through peak and trough. The results of our technique are encouraging. We achieved an annualized return of 14.87% in USD for the 1min sampled dataset ($\lambda$=1). For the 5min and 60min sampled datasets, the annualized returns were 15.49% and 5.00%, respectively, which greatly exceed the Distance Method (DM) in all experimented cases.
Figure 6: The profitability of Full-Cycle Market between 15-Jan-2021 and 01-Oct-2022
We also conduct experiments on conditions where investors have higher or lower tolerance for portfolio volatility. We observe a clear relationship: the lower the $\lambda$, the higher the profitability. This relationship meets our expectation that investors receive a risk premium in exchange for a more volatile portfolio.
Table 6: Annualized Profitability Results in Full-Cycle Market
Interval BH DM OTTλ=0.5 OTTλ=1 OTTλ=2
1 min (TC 0.1%) 12.91% 5.33% 35.54% 14.87% 7.44%
1 min (TC 0.0%) 13.14% 0.73% 79.98% 40.03% 19.97%
5 min (TC 0.1%) 12.91% 0.60% 37.74% 15.49% 7.74%
5 min (TC 0.0%) 13.14% 4.08% 90.03% 44.98% 22.49%
60 min (TC 0.1%) 12.91% 3.93% 9.93% 5.00% 2.51%
60 min (TC 0.0%) 13.14% 5.54% 37.43% 18.96% 9.49%
  • * The arbitrage period is 15-Jan-2021 to 01-Oct-2022.
  • * Participating assets are USD, CAD, GBP, and EUR against ETH.
  • * The annualized returns are stated in percentages.
  • * BH(Buy-Hold), DM(Distance Method), TC(Transaction Cost), OTT(Optimal Trading Technique)
  • * open_threshold and close_threshold are tuned to the most lucrative combination (Table 4).
Another comparison includes transaction costs. We can see that even a 0.1% transaction cost significantly impacts our returns. On the surface, transaction cost charges a flat percentage of trading capital and influences the optimization process’s expectation. Our technique trades on a frequent basis; hence we can see the profitability is almost halved in three examined $\lambda$ values after including transaction costs. The Distance Method, on the other hand, does not have a significant impact due to transaction cost, because it trades less frequently (Table 10).
@@ 5.4.2 Bull Market
Figure 7: The profitability of Bull Market between 01-Jan-2021 and 01-Jan-2022
Table 7 presents the trading results for 2021. A significant growth trend is evident from the Buy-Hold strategy, which yielded a return of 412.09%. Such a circumstance is rare and occurs only in exceptional cases. Compared to full-cycle market trading, both the Distance Method (DM) and our Optimal Trading Technique (OTT)yielded more profit in the bull market. For a standard scenario of $\lambda=1$ and $TC=0.1\%$ over a five-minute period, the OTT achieved an annual return of 36.75% (compared to 19.89% for the DM under the same conditions). Single-sided markets are easier to estimate, and the expected_return and expected_risk are therefore closer to reality. The profitability trend is consistent with that of the full-cycle market. Arbitrages that do not entail transaction fees generate better returns than those with transaction fees. A lower value of $\lambda$ implies a higher acceptance of risk, resulting in a better overall arbitrage return.
Table 7: Annualized Profitability Results in Bull Market
Interval BH DM OTTλ=0.5 OTTλ=1 OTTλ=2
@@
60 min (TC 0.0%) 412.90% 7.46% 45.25% 22.75% 11.37%
  • * The arbitrage period is 01-Jan-2021 to 01-Jan-2022.
  • * Participating assets are USD, CAD, GBP, and EUR against ETH.
  • * The annualized returns are stated in percentages.
  • * BH(Buy-Hold), DM(Distance Method), TC(Transaction Cost), OTT(Optimal Trading Technique)
  • * open_threshold and close_threshold are tuned to the most lucrative combination (Table 4).
@@ 5.4.3 Bear Market
We also have an interest in examining how well our technique performs in a bear market. Our optimization process calculates weights to avoid holding ETH in our portfolio. Therefore in theory, the process should perform as well as in a bull market. Therefore, we experiment with the technique under the bear market in 2022, from 01-Jan-2022 to 31-Dec-2022. During this period, the price of ETH dropped by -61.77%. We examine the profitability with data for this price dip in Table 8.
Figure 8: The profitability of Bear Market between 01-Jan-2022 and 01-Jan-2023
During the bear market in 2022, ETH witnessed a significant drop in all major currencies. The traditional Buy-Hold strategy and most non-market-neutral techniques would experience a loss from the bear market. By not holding ETH at all, both the Distance Method and our Optimal Trading Technique (OTT)achieved positive returns in all experimented cases. We can see from Table 8 that the patterns we found in a full-cycle market also apply here.
Table 8: Annualized Profitability Results in Bear Market
Interval BH DM OTTλ=0.5 OTTλ=1 OTTλ=2
@@
60 min (TC 0.0%) -61.77% 5.53% 33.08% 13.93% 8.52%
  • * The arbitrage period is 01-Jan-2022 to 01-Jan-2023.
  • * Participating assets are USD, CAD, GBP, and EUR against ETH.
  • * The annualized returns are stated in percentages.
  • * BH(Buy-Hold), DM(Distance Method), TC(Transaction Cost), OTT(Optimal Trading Technique)
  • * open_threshold and close_threshold are tuned to the most lucrative combination (Table 4).